Properties of a Family of Partitioned 3D Meshes

David R. O’Hallaron and Jonathan Richard Shewchuk August 21, 1996

CMU-CS-96-141

School of Computer Science Carnegie Mellon University Pittsburgh, PA 15213

Abstract

abstract goes here.

Supported in part by the National Science Foundation under Grant CMS-9318163, by the Advanced Research Projects Agency/CSTO monitored by SPAWAR under contract N00039-93-C-0152, and by a grant from the Intel Corporation. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the of�cial policies, either express or implied, of the U.S. government.

Keywords:

Contents List of Figures

1 Scalability on the T3D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

About this Report

An electronic copy of this report, and the software described herein, can be obtained through the Web page http://www.cs.cmu.edu/�quake/xxx.html. iv

1 Introduction

Figure 1: Scalability on the T3D

[?]